OpenStudy (anonymous):

what is the answer to this system of equations I tried elimination and substitution but it didn't work. Need help. 2x-y+3z=8 x-6y-z=0 -6x+3y-9z=24

OpenStudy (abb0t):

Show the work that you did using the method you feel more strongly in to see where you may have messed up or where you get stuck.

OpenStudy (anonymous):

\[2x-y+3z=8[1]\] \[x-6y-z=0[2]\] \[-6x+3y-9z=24[3]\] \[[2]\] \[x=6y+z\] \[[1]\] \[2(6y+z)-y+3z=8\] \[12y+6z-y+3z=8\] \[11y+9z=8[4]\] \[[3]\] \[-6(6y+z)+3y-9z=24\] \[-36y-6z+3y-9z=24\] \[-33y-15z=24[5]\] \[[4]\times 3\] \[3(11y+9z)=3(8)\] \[33y+27z=24[6]\] \[[5]+[6]\] \[(-33y-15z)+(33y+27z)=(24)+(24)\] \[-15z+27z=48\] \[12z=48\] \[z=?\]

Directrix (directrix):

The common intersection of the system is not a unique point.